# Omnidirectional Multi-Object Tracking

### v1

@@ 5.1 Experiment Setup
Track-By- Detection (TBD) SORT [7] YOLO11 [66] (baseline) 25.83 29.56 0.915 31.02 27.62 24.51
OmniTrackDet (ours) 26.34 (+0.51) 31.11 (+1.55) 0.907 (-0.008) 34.21 (+3.19) 30.52 (+2.90) 22.96 (-1.55)
OmniTrackDA (ours) 29.44 (+3.10) 33.27 (+2.16) 0.927(+0.020) 33.44 (-0.77) 35.16 (+4.64) 25.06 (+2.10)
ByteTrack [81] YOLO11 [66] (baseline) 27.85 32.20 0.896 34.46 31.49 25.15
OmniTrackDet(ours) 28.14 (+0.29) 32.97 (+0.77) 0.870 (-0.026) 37.36 (+2.90) 32.94 (+1.45) 24.29 (-0.86)
OmniTrackDA (ours) 29.58 (+1.44) 34.54 (+1.57) 0.859 (-0.011) 38.14 (+0.78) 34.71 (+1.77) 25.49 (+1.20)
OC-SORT [9] YOLO11 [66] (baseline) 29.26 33.69 0.874 34.22 31.81 27.48
OmniTrackDet (ours) 29.43 (+0.17) 34.11 (+0.42) 0.851 (-0.023) 38.72 (+4.50) 34.48 (+2.67) 25.39 (-2.09)
OmniTrackDA (ours) 30.65 (+1.22) 34.83 (+0.72) 0.838 (-0.013) 36.37 (-2.35) 35.58 (+1.10) 26.76 (+1.37)
HybridSORT [75] YOLO11 [66] (baseline) 29.71 34.16 0.877 34.71 31.70 28.39
OmniTrackDet (ours) 30.00 (+0.29) 34.09 (-0.07) 0.853 (-0.024) 32.32 (-2.39) 35.02 (+3.32) 26.09 (-2.30)
OmniTrackDA (ours) 31.05 (+1.05) 36.06 (+1.97) 0.850 (-0.003) 38.13 (+5.81) 35.08 (+0.06) 27.78 (+1.69)
E2E TrackFormer [53] n.a. 22.22 23.38 0.959 23.83 30.30 16.93
MOTR [79] n.a. 19.78 23.25 0.928 25.44 25.51 15.61
MOTRv2 [82] n.a. 24.68 25.49 0.911 17.05 26.83 22.97
OmniTrackE2E (ours) n.a. 25.12 27.42 0.925 34.99 33.35 19.17
Table 4: Results on the JRDB validation set [51]. The first four groups compare methods under the TBD paradigm, whereas the last group presents a comparison under the E2E paradigm. In the TBD paradigm, each method is evaluated under three detection methods: the baseline with YOLO11 [66] as the detector, the OmniTrackDet detector, and OmniTrackDA. The numbers represent the improvement relative to the previous line’s method.
@@
We utilize the CLEAR metrics [6], which include MOTA, False Positives (FP), False Negatives (FN), and others, in conjunction with IDF1 [60], OPSA [51], and HOTA [49] to provide a comprehensive evaluation of tracking performance across various dimensions. MOTA emphasizes detector performance, while IDF1 evaluates the tracker’s ability to maintain consistent identities. In contrast, HOTA integrates association accuracy, and localization accuracy, establishing itself as an increasingly significant metric for assessing tracking algorithms.
@@ 5.2 Comparison with State of the Art
In Tab. 3, we present a comparison between OmniTrack and state-of-the-art methods on the QuadTrack test set. This dataset is particularly challenging, characterized by a wide FoV and rapid, non-linear sensor motion, which introduces significant complexities for traditional MOT algorithms. Despite these challenges, our method outperforms existing approaches, achieving the highest HOTA scores: $19.87\%$ for the E2E group and $23.45\%$ for the TBD group.
@@ 5.3 Paradigm Comparison
In the same tracking method, OmniTrackDet consistently outperforms YOLO11-X [66], showing an average improvement of $0.2\%$ in HOTA and $0.6\%$ in IDF1. This highlights the superior performance of OmniTrackDet over YOLO11-X [66]. Furthermore, when comparing OmniTrackDet to OmniTrackDA, the latter shows an average increase of $1.7\%$ in HOTA and $1.4\%$ in IDF1, demonstrating the effectiveness of the feedback mechanism (Sec. 3.1). In the E2E paradigm, OmniTrackE2E achieved the best result HOTA of $25.12\%$ and IDF1 of $27.42\%$.
@@ 5.4 Ablation Study
In Tab. 6, we evaluate the effectiveness of DynamicSSM in the CircularStatE, comparing it with other common designs such as Conv and MLP. The results from experiments ➁, ➂, and ➃ demonstrate a clear advantage for DynamicSSM. Experiments ➄, ➅, and ➆ further show that applying DynamicSSM to $S_{4}$ yields the best performance. This is because $S_{4}$ contains both semantic and texture information, which makes it easier for DynamicSSM to extract distortion and color details, thus aiding in calibration.
@@ 6 Conclusion
This paper presents OmniTrack, a multi-object tracking framework tailored for panoramic images, effectively addressing key challenges like geometric distortion, low resolution, and lighting inconsistencies. Central to OmniTrack is a feedback mechanism that reduces uncertainty in large-FoV tracking. The framework incorporates Tracklets Management for temporal stability, FlexiTrack Instance for rapid localization and association, and the CircularStatE Module to mitigate distortion and improve visual consistency. Additionally, we present QuadTrack, a cross-campus multi-object tracking dataset collected using a quadruped robot to support dynamic motion scenarios. This challenging dataset is designed to advance research in omnidirectional perception for robotics. Experiments verify that OmniTrack achieves state-of-the-art performance on public JRDB and the established QuadTrack datasets, demonstrating its effectiveness in handling panoramic tracking tasks.

### v2

@@ 5.1 Experiment Setup
Track-By- Detection (TBD) SORT [7] YOLO11 [66] (baseline) 25.83 29.56 0.915 31.02 27.62 24.51 49.18
OmniTrackDet (ours) 26.34 (+0.51) 31.11 (+1.55) 0.907 (-0.008) 34.21 (+3.19) 30.52 (+2.90) 22.96 (-1.55) 12.14
OmniTrackDA (ours) 29.44 (+3.10) 33.27 (+2.16) 0.927(+0.020) 33.44 (-0.77) 35.16 (+4.64) 25.06 (+2.10) 11.78
ByteTrack [81] YOLO11 [66] (baseline) 27.85 32.20 0.896 34.46 31.49 25.15 50.36
OmniTrackDet(ours) 28.14 (+0.29) 32.97 (+0.77) 0.870 (-0.026) 37.36 (+2.90) 32.94 (+1.45) 24.29 (-0.86) 12.24
OmniTrackDA (ours) 29.58 (+1.44) 34.54 (+1.57) 0.859 (-0.011) 38.14 (+0.78) 34.71 (+1.77) 25.49 (+1.20) 11.83
OC-SORT [9] YOLO11 [66] (baseline) 29.26 33.69 0.874 34.22 31.81 27.48 46.33
OmniTrackDet (ours) 29.43 (+0.17) 34.11 (+0.42) 0.851 (-0.023) 38.72 (+4.50) 34.48 (+2.67) 25.39 (-2.09) 11.59
OmniTrackDA (ours) 30.65 (+1.22) 34.83 (+0.72) 0.838 (-0.013) 36.37 (-2.35) 35.58 (+1.10) 26.76 (+1.37) 11.13
HybridSORT [75] YOLO11 [66] (baseline) 29.71 34.16 0.877 34.71 31.70 28.39 44.34
OmniTrackDet (ours) 30.00 (+0.29) 34.09 (-0.07) 0.853 (-0.024) 32.32 (-2.39) 35.02 (+3.32) 26.09 (-2.30) 11.65
OmniTrackDA (ours) 31.05 (+1.05) 36.06 (+1.97) 0.850 (-0.003) 38.13 (+5.81) 35.08 (+0.06) 27.78 (+1.69) 10.96
E2E TrackFormer [53] n.a. 22.22 23.38 0.959 23.83 30.30 16.93 7.38
MOTR [79] n.a. 19.78 23.25 0.928 25.44 25.51 15.61 12.73
MOTRv2 [82] n.a. 24.68 25.49 0.911 17.05 26.83 22.97 13.01
OmniTrackE2E (ours) n.a. 25.12 27.42 0.925 34.99 33.35 19.17 11.64
Table 4: Results on the JRDB validation set [51]. The first four groups compare methods under the TBD paradigm, whereas the last group presents a comparison under the E2E paradigm. In the TBD paradigm, each method is evaluated under three detection methods: the baseline with YOLO11 [66] as the detector, the OmniTrackDet detector, and OmniTrackDA. The numbers represent the improvement relative to the previous line’s method. The FPS metric is measured on a single RTX 3090 GPU with an image resolution of $4160\times 480$.
@@
We employ the CLEAR metrics [6], including MOTA, DetA, and AssA, alongside IDF1 [60], OPSA [51], and HOTA [49] for a comprehensive tracking performance evaluation. MOTA is primarily influenced by detector performance, IDF1 measures identity preservation, and HOTA integrates association and localization accuracy, making it increasingly pivotal for tracking assessment.
@@ 5.2 Comparison with State of the Art
In Tab. 3, we present a comparison between OmniTrack and state-of-the-art methods on the QuadTrack test set. This dataset is particularly challenging, characterized by a panoramic FoV and rapid, non-linear sensor motion, which introduces significant complexities for traditional MOT algorithms. Despite these challenges, our method outperforms existing approaches, achieving the highest HOTA scores: $19.87\%$ for the E2E group and $23.45\%$ for the TBD group.
@@ 5.3 Paradigm Comparison
In the TBD method, OmniTrackDet consistently outperforms YOLO11-X [66], showing an average improvement of $0.2\%$ in HOTA and $0.6\%$ in IDF1. Despite OmniTrackDet not having a speed advantage, it achieves notable improvements in accuracy. Furthermore, when comparing OmniTrackDet to OmniTrackDA, the latter shows an average increase of $1.7\%$ in HOTA and $1.4\%$ in IDF1, demonstrating the effectiveness of the feedback mechanism. In the E2E paradigm, OmniTrackE2E achieved the best result HOTA of $25.12\%$ and IDF1 of $27.42\%$.
@@ 5.4 Ablation Study
In Tab. 6, we evaluate the effectiveness of DynamicSSM in the CircularStatE, comparing it with other common designs such as Conv and MLP. The results from experiments ➁, ➂, and ➃ demonstrate a clear advantage for DynamicSSM. Experiments ➄, ➅, and ➆ further show that applying DynamicSSM to $S_{4}$ yields the best performance. where $S_{5}$, $S_{4}$, and $S_{3}$ impact MOT results. Since $S_{4}$ contains both high-level semantic and low-level geometric features, its effect is the most pronounced.
@@ 6 Conclusion
This paper presents OmniTrack, a multi-object tracking framework tailored for panoramic images, effectively addressing key challenges like geometric distortion, low resolution, and lighting inconsistencies. Central to OmniTrack is a feedback mechanism that reduces uncertainty in panoramic-FoV tracking. The framework incorporates Tracklets Management for temporal stability, FlexiTrack Instance for rapid localization and association, and the CircularStatE Module to mitigate distortion and improve visual consistency. Additionally, we present QuadTrack, a cross-campus multi-object tracking dataset collected using a quadruped robot to support dynamic motion scenarios. This challenging dataset is designed to advance research in omnidirectional perception for robotics. Experiments verify that OmniTrack achieves state-of-the-art performance on public JRDB and the established QuadTrack datasets, demonstrating its effectiveness in handling panoramic tracking tasks.
@@
Acknowledgment
This work was supported in part by the National Natural Science Foundation of China (No. 62473139 and No. 12174341), in part by Zhejiang Provincial Natural Science Foundation of China (Grant No. LZ24F050003), and in part by Shanghai SUPREMIND Technology Co. Ltd.